OpenStudy (anonymous):

how to solve algebraically \[10 = \frac{2 times X}{2 + X}\] is it possible?

OpenStudy (anonymous):

I'll try \[10 = \frac{ 2 }{ 2+ x } \times \frac{ x }{ 1 }\] not sure if that helped

OpenStudy (anonymous):

multiplay by (2+x) on both sides \[10(2+x) = 2 \times x\] solve \[20+10x = 2 \times x\] and that is? \[8x = - 20\]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

cool :D didn't think I could do anything to this